### v1

# Scalable Hybrid quantum Monte Carlo simulation of U(1) gauge field coupled to fermions on GPU

Abstract
We develop a GPU-accelerated hybrid quantum Monte Carlo (QMC) algorithm to solve the fundamental yet difficult problem of $U(1)$ gauge field coupled to fermions, which gives rise to a $U(1)$ Dirac spin liquid state under the description of (2+1)d quantum electrodynamics QED3. The algorithm renders a good acceptance rate and, more importantly, nearly linear space-time volume scaling in computational complexity $O(N_{\tau}V_{s})$, where $N_{\tau}$ is the imaginary time dimension and $V_{s}$ is spatial volume, which is much more efficient than determinant QMC with scaling behavior of $O(N_{\tau}V_{s}^{3})$. Such acceleration is achieved via a collection of technical improvements, including (i) the design of the efficient problem-specific preconditioner, (ii) customized CUDA kernel for matrix-vector multiplication, and (iii) CUDA Graph implementation on the GPU. These advances allow us to simulate the $U(1)$ Dirac spin liquid state with unprecedentedly large system sizes, which is up to $N_{\tau}\times L\times L=660\times 66\times 66$, and reveal its novel properties. With these technical improvements, we see the asymptotic convergence in the scaling dimensions of various fermion bilinear operators and the conserved current operator when approaching the thermodynamic limit. The scaling dimensions find good agreement with field-theoretical expectation, which provides supporting evidence for the conformal nature of the $U(1)$ Dirac spin liquid state in the QED3. Our technical advancements open an avenue to study the Dirac spin liquid state and its transition towards symmetry-breaking phases at larger system sizes and with less computational burden.
I Introduction
I.1 General Background
The emergence of Dirac spin liquid (DSL), i.e. $U(1)$ gauge field coupled to gapless relativistic spinons at (2+1)d [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12], or, in high-energy physics terms, a deconfined phase in 3D quantum electrodynamics (QED3) coupled to $N_{f}$ flavor massless fermions [13, 14, 15, 16, 17, 18, 19, 20], has hold the attention of condensed matter and high-energy physicists for decades. QED3 is expected to realize a conformal field theory (CFT) for large $N_{f}$ [19, 17], which contains many interesting scaling operators including fermion bilinears operators in the form of $SU(N)$ flavor currents and various fermion bilinear terms, as well as monopole operators, which generate $U(1)$ magnetic fluxes and transform non-trivially on lattice.
Over the years, the QED3 CFT becomes the prototypical problem for testing perturbative and non-perturbative methods to study strongly coupled conformal field theory in 2+1 dimensions in the high energy community [13, 14, 15, 16, 19, 18, 17, 20]. In condensed matter physics, QED3 state has been argued to arise as a ground state of certain frustrated quantum spin models with Heisenberg-type Hamiltonian, with no magnetic order (no spontaneous symmetry breaking) and yet the emergent spinon exhibiting Dirac dispersion and their interactions mediated by the fluctuating $U(1)$ gauge field. It is in this context such a gapless and critical state is known as $U(1)$ DSL [21, 1, 2, 3, 4, 5, 22, 23, 24, 25, 26, 27, 28, 29, 6, 7, 8, 9, 30, 31, 32, 10, 11, 12, 33, 34].
From a field-theoretical perspective, the basic properties of QED3 are understood, e.g. one knows the primary fields and their roughly estimated scaling dimensions [18, 17, 20]. However, concrete lattice model calculations, with the precise properties of multi-point correlation functions and their evolution across phase transition to other symmetry-breaking phases, tuned by quantum and thermal knobs such as the fluctuation strength, magnetic field, temperature, etc, are yet unknown. This, to the authors viewpoint, is mainly due to the lack of controlled computational scheme that can solve this strongly interacting theory in an unbiased manner and extrapolate, for example, the correlation functions of various fermion bilinears, from finite space-time volume to the thermodynamic limit. It is such difficulty that motivates us to develop the scalable hybrid quantum Monte Carlo on GPU to solve this problem in this work.
At the condensed matter experimental front, the signature of (proximate) $U(1)$ DSL has been actively reported in frustrated quantum magnets in recent years, in particular triangular lattice antiferromagnets YbZn2GaO5 [35, 36] and the $A$-YbSe2 delafossites [37, 38] and kagomé antiferromagnets [39, 40, 41, 42, 43, 44, 45]. For example, in the recently synthesized kagomé material YCu3(OD)6Br2[Br1-x(OD)x], thermodynamic measurements have revealed clear $T^{2}$ dependence in specific heat, suggesting gapless degrees of freedom with linear dispersion in this electronically gapped Mott insulator [39]. Subsequently, inelastic neutron scattering reveals spin spectral signature of the Dirac cone continuum for a nonmagnetic ordered state at low temperature [40, 41], and magnetic torqe measurements found unconventional quantum oscillations near the 1/9 magnetization plateau in this Mott insulator [45], all these results are pointing towards a DSL state in the material.
In light of these progresses, it is timely to develop a quantum many-body computation scheme that can access the QED3 $U(1)$ DSL state with ease and control, such that one can address and understand the theoretical questions and make ossible connections to experimental observations.
I.2 Computational Challenges
As mentioned above, theoretical approaches cannot arrive at controlled results unless $N_{f}$ is large, but the physical situation is at $N_{f}=2$ for spin-1/2 fermions. Considering the fermion doubling on the lattice, there will be $2N_{f}$ Dirac fermions and the lattice symmetry is enlarged to an emergent $SU(2N_{f})=SU(4)$ symmetry. In terms of concrete lattice model calculations, only very recently, unbiased numerical results of QED3 on finite size lattices from determinant quantum Monte Carlo (DQMC) simulations start to appear [6, 7, 9], but the simulation can not access sufficiently large system sizes and low temperature as it suffers heavily from unfriendly computational complexity of fermion determinant and slow Monte Carlo dynamics due to local update in the large configurational space spanned by the continuous $U(1)$ variables in (2+1)d space-time lattice.
To be more specific, even with specially designed fast updates (it is still a local update in the gauge field configuration) [6], the computational complexity of DQMC remains challenging at $O(N_{\tau}V_{s}^{3})$, with $N_{\tau}$ the imaginary time dimension and $V_{s}=L^{2}$ the number of spatial sites ($L$ is the linear dimension). The primary source of this scaling stems from the explicit determinant evaluation in the partition function. Calculating the acceptance ratio in DQMC depends on the equal-time Green’s function, and updating the Green’s function after sweeping through the space-time lattice renders a complexity of either $O(N_{\tau}V_{s}^{3})$ (with fast updates) or $O(N_{\tau}V_{s}^{4})$ (without fast updates).
Moreover, the $U(1)$ DSL phase is a critical phase characterized by gapless dynamics of both the gauge field and the fermion. In such phase, the simulation usually experiences critical slowing down, i.e. the Monte Carlo autocorrelation time $\tau_{\text{int}}$ of physical observables grows with system size as $\tau_{\text{int}}\sim L^{z_{\text{MC}}}$, where $z_{\text{MC}}$ is the algorithmic dynamical exponent reflecting the efficiency of the Monte Carlo update scheme. DQMC with Metropolis update for critical phases often renders a larger $z_{\text{MC}}$ than its classical counterpart ($z_{\text{MC}}=2.17$ for 2D classical Ising model at its critical point) [46, 47]. This inefficiency becomes severe for the continuous gauge field with highly anisotropic terms in the Lagrangian in Eqs. (1), (2) and (3), as the ratio between temporal and spatial terms is of order $\frac{1}{\Delta\tau^{2}}$, with $\Delta\tau$ being the Trotter step size. The dynamics of local update is thus dominated by the temporal term in Lagrangian. More specifically, local updates of the form $\phi^{\prime}_{ij,\tau}=\phi_{ij,\tau}+\Delta\phi$ suffer from low acceptance rates when the gauge coupling $J$ is small—typically in the $U(1)$ DSL phase. Since the energy difference of such updates is inversely proportional to $J$, a smaller $J$ requires a smaller $\Delta\phi$ update to maintain a reasonable acceptance ratio. And a smaller $\Delta\phi$ in turn results in larger autocorrelation time in the sampled configurations.
Therefore the DQMC seems difficult to handle the problem, whether there are other Monte Carlo methods that could come to help? The answer is yes and it is the hybrid quantum Monte Carlo (HQMC) method [48, 49, 50, 51, 52, 53]. At least on paper, it enjoys two clear advantages over DQMC in studying this problem: (i) instead of the local update in DQMC, the Hamilton’s equation of motion in HQMC updates the gauge field configurations globally, while preserving a good acceptance rate; (ii) with the help of psudo-fermion transformation, the evaluation of the fermionic determinant, which has $O(N_{\tau}V_{s}^{3})$ scaling, is replaced by solving a linear system with iterative methods, which can have as low a computation complexity as $O(N_{\tau}V_{s})$, and is also amenable to parallel computing like GPU acceleration. So, in principle, HQMC is a much more efficient method than DQMC.
However, actual application of the HQMC in condensed matter models faces many challenges, which are often problem specific. Previous attempts of employing HQMC on the Hubbard model did not show significant advantages  [48, 54]. In particular, in a recent attempt [53], while the ergodicity issue is successfully solved by a compactification method, the overall time complexity of the HQMC algorithm is not much better than conventional DQMC, because of the diverging iteration number complexity of the linear equation solver. Therefore, doubts are cast upon the applicability of HQMC on strongly correlated systems in condensed matter physics.
But the more recent developments are very encouraging. The applications of HQMC to the spin-fermion models for the investigation of the non-Fermi-liquid and quantum critical metallic state [55, 56], apparently enjoy successes, both in the physical discoveries and in computation technologies. In the spin-fermion models, the bosonic fields originate from the physical spin degrees of freedom which provide the critical fluctuations [57, 58, 59], and are different from the auxiliary bosonic field in Hubbard model which comes from the Hubbard-Stratonovich transformation while decoupling the density-density interaction. It is found that by applying the HQMC on spin-fermion models with various optimization techniques e.g. preconditioned conjugate gradient (pCG) method, matrix-free multiplication and GPU parallel computing, the authors of Ref. [55, 56] significantly scale up the simulation size, which is up to $N_{\tau}\times L\times L=200\times 80\times 80$. The overall computational complexity is also reduced to be linear with respect to time $N_{\tau}$ and spatial volume $V_{s}$, i.e. $O(N_{\tau}V_{s})$.
I.3 Our GPU-accelerated HQMC solution
Inspired by such development. In this paper, we apply HQMC to simulate a $U(1)$ gauge field model coupled to $N_{f}=2$ fermions at (2+1)d space-time on GPU, and manage to scale the simulation size up to $N_{\tau}\times L\times L=660\times 66\times 66$, which, to the best of our knowledge, is a huge improvement in simulation size compared to all previous DQMC studies of this model [6, 7, 9], which is typically up to $N_{\tau}\times L\times L=200\times 20\times 20$, as well as three times larger than the HQMC study of a different spin-fermion model in Ref. [55, 56]. Such improvements are achieved by the implementation of the following several GPU related optimization techniques tailored to this problem, including
  1. 1.
    the problem-specific preconditioner computation,
  2. 2.
    customized CUDA kernel for matrix-vector multiplication that exploits the specific sparsity pattern, and
  3. 3.
    CUDA graph acceleration that minimizes kernel launch overhead.
With these advances, in this fundamental model, computational complexity is now successfully reduced from DQMC’s $O(N_{\tau}V_{s}^{3})$ to a linear complexity $O(N_{\tau}V_{s})$ with respect to imaginary time $N_{\tau}$ and spatial volume $V_{s}$.
Using such GPU-accelerated HQMC, we are able to investigate the existence of $U(1)$ DSL state in QED3 model with ease and control. We study the correlation function of different fermion bilinear operators in DSL state such as spin correlation functions and bond correlations, and find they have the same scaling dimension at the thermodynamic limit, as they belong to the fermion bilinear terms (adjoints) of the emergent $SU(4)$ group and share the same scaling dimension of $\Delta_{\rm adj}\in(1.4,1.7)$ [3, 18]. We reveal such a property for the case of $N_{f}=2$ ($N=4$ due to fermion doubling in DSL phase) QED3, by first benchmarking with DQMC at small system sizes and then performing large-scale simulation up to to $660\times 66\times 66$. We also observe that the flux-flux correlations is dictated by conserved current scaling dimension $\Delta_{J}=2$ of the $SU(4)$ group, such that the long-time correlations of the local flux approach $1/\tau^{4}$ asymptotically. We believe this work opens an avenue to explore the generic phase diagram of the model and the novel phase transitions between $U(1)$ DSL and other phases, in a computationally efficient and scalable manner.
II Models and Methods
II.1 Models
Figure 1: Lattice model of QED3. Fermions live on the sites of the lattice represented by blue dots, while gauge fields live on the bonds of the lattice represented by yellow dots. Fermion hops between nearest-neighbor sites with a phase $e^{i\phi_{ij}}$. The plot shows two adjacent square lattice layers, which are connected with temporal bonds represented as black dashed line. Gauge fields on these temporal bonds are set to be zero. Within a square lattice layer, the four bond colors represent the four families of bonds, which will be used in the checkerboard decomposition, as shown in Eq. (19); within each family, the fermion hopping terms represented by the bonds all mutually commute. In the bottom right cube, the circles on the facets denote the positive directions of the magnetic flux going out of the cube.
The models we study are quantum electrodynamics (QED3) defined on a cubic space-time lattice (square spatial lattice), coupled with two-flavor fermions, as shown in Fig. 1 [6, 7]. Here, we start with the action of these lattice QED3 models: $S=S_{F}+S_{B}=\int_{0}^{\beta}d\tau\left(L_{F}+L_{B}\right)$, where $\beta$ is the inverse temperature and $\tau$ is the imaginary time. The first term
$L_{F}=$ $\sum_{\langle ij\rangle,\alpha}\psi_{i,\tau,\alpha}^{\dagger}\left(\partial_{\tau}\delta_{ij}-te^{i\phi_{ij,\tau}}\right)\psi_{j,\tau,\alpha}+\text{ h.c. }$ (1)
is the fermionic Lagrangian, where $\psi_{i,\tau,\alpha}$ is the fermionic field with spatial index i[1,L]i\ $\in$[1,L], temporal index $\tau\in[1,N_{\tau}]$ and flavor index $\alpha\in\left\{\uparrow,\downarrow\right\}$. $t=1$ is the hopping amplitude and set as the energy unit throughout the paper. $\phi_{ij,\tau}$ is a bosonic scalar that lives on a bond $ij$, and acts as a $U(1)$ gauge field. The dynamics of the $U(1)$ gauge field is described by the bosonic action $S_{B}$.
In this paper, we consider two $U(1)$ gauge field Lagrangians, namely compact $U(1)$ gauge field
$L_{B}^{\textmd{c}}=\frac{1}{J\Delta\tau^{2}}\sum_{\langle ij\rangle}\left[1-\cos\left(\phi_{ij,\tau+1}-\phi_{ij,\tau}\right)\right]+K\sum_{\square}\cos(\operatorname{curl}\phi),$ (2)
and non-compact $U(1)$ gauge field
$L_{B}^{\textmd{nc}}=\frac{1}{J\Delta\tau^{2}}\sum_{\langle ij\rangle}\left[\phi_{ij,\tau+1}-\phi_{ij,\tau}\right]^{2}+K\sum_{\square}\cos(\operatorname{curl}\phi),$ (3)
where $J$ and $K$ are coupling constants, and $\Delta\tau=0.1$ is the imaginary time step introduced in the Trotter decomposition [6] in the lattice Monte Carlo simulation. The discretized version of $\operatorname{curl}\phi$ for a spatial plaquette $\square$ (which is also seen as magnetic flux through the plaquette) is defined as $\theta^{xy}_{\square}=\operatorname{curl}\phi=\sum_{ij\in\textmd{edge}(\square)}(-1)^{ij}\phi_{ij}$, where $(-1)^{ij}$ is $1$ if $ij$ is along the counterclockwise direction of the plaquette and 1-1 if opposite [60, 13], as illustrated in the bottom right cube in Fig. 1. Similarly, in the cubic space-time lattice, the curl of a temporal plaquette, i.e. $\theta^{\tau x}$ or $\theta^{\tau y}$, can also be defined. If we fix the gauge field of temporal bonds to be zero, then the curl of a temporal plaquette is $\pm(\phi_{ij,\tau+1}-\phi_{ij,\tau})$, which consists of the $\tau$-derivative term in the gauge field Lagrangians. It is then clear that $L_{B}^{\textmd{c}}$ has $2\pi$-periodicity in the magnetic flux: $\theta\to\theta+2n\pi,n\in\mathbb{Z}$. Thus, $L_{B}^{\textmd{c}}$ describes a compact $U(1)$ gauge field theory in the sense that the magnetic flux is physical only in the range of $(-\pi,\pi]$, while the other values of $\theta$ are unphysical and need to be wound into the $(-\pi,\pi]$ interval. Due to $2\pi$-periodicity (the fermionic Lagrangian $L_{F}$ also has $2\pi$-periodicity by definition [61]), the compact QED3 is known to host magnetic monopoles, which contains distinct physics from a non-compact theory e.g. $L^{\textmd{nc}}_{B}$ in Eq. (3), where the first term breaks the $2\pi$-periodicity. For example, the proliferation of magnetic monopoles in compact QED3 will lead to the emergence of a mass gap in the gauge field, through the dual superconductor mechanism [62, 63, 19], while the non-compact QED3 is generally believed to remain deconfined. We therefore focus on the results on non-compact QED3 in the main text and leave the compact QED3 results in the Supplemental Materials (SM) [64]. Although from the algorithmic point of view, there is no difference in their implementation and performance of the corresponding simulations.
In our numerical simulation, the quadratic fermion $\psi$ can be traced out, yielding
$\operatorname{Tr}_{\psi}\left[e^{-S_{f}}\right]=\prod_{\alpha}\left[\operatorname{det}\left(\mathbf{1}+\prod_{\tau=1}^{N_{\tau}}B_{\tau,\alpha}\right)\right]=\prod_{\alpha}\operatorname{det}M_{\alpha},$ (4)
where $B_{\tau,\alpha}=e^{V_{\tau,\alpha}}$, and $V_{\tau,\alpha}$ is the exponential of fermion-gauge coupling matrix, whose elements are determined by gauge field $\Delta_{\tau}e^{i\phi_{ij,\tau}}$. The matrix
$M_{\alpha}$ $=\begin{pmatrix}\mathbf{1}&0&0&\cdots&0&B_{N_{\tau},\alpha}\\ -B_{1,\alpha}&\mathbf{1}&0&\cdots&0&0\\ 0&-B_{2,\alpha}&\mathbf{1}&\cdots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\cdots&\mathbf{1}&0\\ 0&0&0&\cdots&-B_{N_{\tau}-1,\alpha}&\mathbf{1}\end{pmatrix}$. (5)
is of size $N_{\tau}V_{s}\times N_{\tau}V_{s}$, where $V_{s}=L\times L$ is the number of spatial lattice sizes and $N_{\tau}$ is the number of imaginary time slices. In our simulation, $N_{\tau}=10L$ ($\beta=\frac{1}{T}=L$ and $\Delta\tau=0.1$) to ensure that the temperature is low enough for a given $L$. Furthermore, it can be shown [6, 65] that $\operatorname{det}M_{\alpha}\in\mathbb{R}$ and $\operatorname{det}M_{\uparrow}=\operatorname{det}M_{\downarrow}$, thus the simulation is free from the sign problem. Based on the formula above, the partition function can then be expressed as
$Z=\int\left[\delta\phi\right]e^{-S_{B}(\phi)}\prod_{\alpha}\operatorname{det}M_{\alpha}(\phi),$ (6)
where $S_{B}(\phi)$ can be either compact or non-compact bosonic gauge action. The Monte Carlo simulation will then be applied to sample the gauge field $\phi$ and explore the Hilbert space defined by the partition function in Eq. (6), which includes the fermion determinant computation for each $\phi$ configuration.
II.2 HQMC algorithm
In this subsection, we will briefly introduce the HQMC algorithm applied to simulate the partition function Eq. (6). This algorithm originates from lattice quantum chromodynamics simulation [51], and later finds its application in a broader class of statistical problems [66, 67]. An application in the Hubbard model setting of condensed matter physics can be found in Ref. [53]. In the most recent progress, the method is significantly accelerated with the help of GPU in the study of spin-fermion models and the associated non-Fermi-liquid and quantum critical metallic states [55, 56].
To sample the configurational space of Eq. (6), but avoiding the cubic complexity in computing $\operatorname{det}M$, a commonly used technique in HQMC is to introduce pseudofermion fields $\eta$ and convert the determinant into the exponential,
$Z=\int\left[\delta\phi\delta\eta\right]e^{-S_{B}(\phi)-\eta^{\dagger}\left[M^{\dagger}(\phi)M(\phi)\right]^{-1}\eta}.$ (7)
Here, the Gaussian integral of the complex fields $\eta\in\mathbb{C}^{V_{s}N_{\tau}}$ has been conversely used, and the identity $\operatorname{det}M_{\uparrow}=\operatorname{det}M_{\downarrow}\in\mathbb{R}$ is assumed, and proved in Ref. [6]. Thus, the fermionic flavor $\alpha$ is omitted.
Next, we introduce the Hamiltonian dynamics, which provides a mechanism to evolve the gauge field continuously along an equal-energy trajectory. Thus the phase space is explored effectively with high acceptance rate. We further add a momentum field $p_{ij,\tau}$ canonical to the gauge field $\phi_{ij,\tau}$:
$Z=\int\left[\delta\phi\delta p\delta\eta\right]e^{-H\left(\phi,p,\eta\right)},$ (8)
where
$H\left(\phi,p,\eta\right):=$ $S_{B}(\phi)+\sum_{ij,\tau}p_{ij,\tau}^{2}+\eta^{\dagger}\left[(M^{\dagger}(\phi)M(\phi)\right])^{-1}\eta.$ (9)
In the partition function Eq. (8), variables $p$ and $R:=\left[M^{\dagger}(\phi)\right]^{-1}\eta$ obey the Gaussian distribution. Then, the HQMC algorithm that samples the partition function Eq. (8) can be readily introduced as the following steps. For a given initial bosonic field configuration $\phi_{0}\in\mathbb{R}^{2V_{s}N_{\tau}}$,
  • (i)
    sample $p_{0}\in\mathbb{R}^{2V_{s}N_{\tau}}$ and $R\in\mathbb{C}^{V_{s}N_{\tau}}$ from the Gaussian distribution; solve for $\eta=M^{\dagger}(\phi_{0})R$;
  • (ii)
    evolve $(\phi_{0},p_{0})$ by the Hamiltonian dynamics described by the differential equations:
    $\dot{p}_{ij,\tau}$ $=-\frac{\partial H}{\partial\phi_{ij,\tau}}$ (10)
    $\dot{\phi}_{ij,\tau}$ $=\frac{\partial H}{\partial p_{ij,\tau}}=2p_{ij,\tau},$ (11)
    which will yield the proposal variables $(\phi_{t},p_{t})$; note that the $t$ here is the time parameter for the differential equation;
  • (iii)
    check $(\phi_{t},p_{t})$ with the Metropolis-Hastings rule to get the acceptance ratio:
    $r_{\textmd{MH}}=\min\left(1,e^{H(\phi_{0},p_{0},\eta)-H\left(\phi_{t},p_{t},\eta\right)}\right);$ (12)
    output $(\phi_{t},p_{t})$ if accepted and $(\phi_{0},p_{0})$ otherwise;
  • (iv)
    set $\phi_{0}$ as the output $\phi$ and go to the next iteration starting from (i).
Some further clarifications of this algorithm are in order.
First, the Hamiltonian dynamics has the property that the phase space volume is conserved, according to the Liouville theorem [53], which ensures that $H(\phi_{0},p_{0},\eta)=H\left(\phi_{t},p_{t},\eta\right)$ and that the proposal variables $(\phi_{t},p_{t})$ are almost always accepted, if the differential equation is solved exactly.
Second, we apply the well-known Leapfrog method [53] to solve the differential equation Eq. (11). It has the property of time reversibility, which, together with the Metropolis-Hasting rule, guarantees that the Monte Carlo updates fulfill the detailed balance condition [68]. The Leapfrog method is described as follows. At the $n$-th step, to evolve a given $(\phi_{n-1},p_{n-1})$ for a time step $\Delta t$,
$p_{n-\frac{1}{2}}$ $=p_{n-1}+F\left(\phi_{n-1}\right)\frac{\Delta t}{2}$
$\phi_{n}$ $=\phi_{n-1}+2p_{n-\frac{1}{2}}\Delta t$ (13)
$p_{n}$ $=p_{n-\frac{1}{2}}+F\left(\phi_{n}\right)\frac{\Delta t}{2}.$
Here, the force $F(\phi):=-\frac{\partial H}{\partial\phi}$ consists of both the bosonic force $\frac{\partial S_{B}}{\partial\phi}$ and the fermionic force $\frac{\partial}{\partial\phi}\left\{\eta^{\dagger}\left[(M^{\dagger}(\phi)M(\phi)\right]^{-1}\eta\right\}$, the derivation of which is shown in SM Sec. SI. Calculation of Fermionic Force for QED3. In this calculation, generically a matrix vector of the form $\left[M^{\dagger}M\right]^{-1}v$ is involved, which is equivalent to solving a linear system $\left[M^{\dagger}M\right]X=v$, and will be calculated using a preconditioned conjugate gradient (pCG) algorithm, explained in Sec. II.4. In this Leapfrog method, the time step size $\Delta t$ and the number of leapfrog steps $N_{\textmd{leapfrog}}$ in a Monte Carlo sweep are two tunable parameters. Larger $N_{\textmd{leapfrog}}$ benefits from larger change in the proposed configuration, but suffers from the larger accumulated error in the iterative process. We choose $N_{\textmd{leapfrog}}=3$ in our implementation. The time step size $\Delta t$, on the other hand, is adaptively tuned to maintain a moderate acceptance rate.
Third, since the update scheme in HQMC is based on solving the Hamiltonian differential equation, it is naturally continuous. As a result, the random walk in HQMC is limited to a continuous region the configuration space. But in the QED3 model studied in this paper, there are two non-continuous regions, namely those with $(\det M_{\uparrow},\det M_{\downarrow})=(+,+)$ and $(\det M_{\uparrow},\det M_{\downarrow})=(-,-)$. They are separated by the gauge configurations where $\det M\approx 0$, which themselves have diminishing contribution in the partition function Eq. (6), but prevents the HQMC update from going from one region to the other. This leads to the ergodicity issue, shared by the application of HQMC on other models [53] as well. One common solution is to introduce another legitimate update that is ergodic. Then the combination of the multiple updates will cure the ergodicity issue in HQMC [66]. In Ref. [53], the solution of this issue is to decouple the Hamiltonian into two channels and then to complexify the configuration space, which is more problem-specific. In the QED3 studied here, we find that in the small $J$ region, which is the $U(1)$ DSL phase region and the focus of this paper, the ergodicity issue is numerically negligible. So we directly apply the HQMC algorithm in this region. The validity is benchmarked by the DQMC algorithm and is shown in Sec. II of SM [64].
II.3 Stochastic estimation of fermionic observables
For a given bosonic configuration $\phi$, the fermionic observables are easily computed by using the standard stochastic estimation procedure [69, 51, 70]. More specifically, the single-particle Green’s function, $G_{ij}=\langle c_{i}(\tau)c^{\dagger}_{j}(0)\rangle$, which is obtained from the inversion of the $M$ matrix, $M^{-1}$. The direct inversion of $M$ is numerically expensive and scales as $O((V_{s}N_{\tau})^{3})$. Alternatively, the standard stochastic procedure to compute $M^{-1}$, as an iterative method, is usually more efficient. In this procedure, first sample a random vector $\xi\in\mathbb{C}^{V_{s}N_{\tau}}$, whose entries are i.i.d. and have the orthogonality property $\mathbb{E}_{\xi}\xi^{*}_{i}\xi_{j}=\delta_{ij}$. Then, $\sum_{l}\left[M^{-1}\right]_{il}\xi_{l}\xi^{*}_{j}$ will give an unbiased estimation of $\left[M^{-1}\right]_{ij}$, since
$\mathbb{E}_{\xi}\sum_{l}\left[M^{-1}\right]_{il}\xi_{l}\xi^{*}_{j}=\left[M^{-1}\right]_{ij}.$
Here, the random vector can be drawn from any distribution that satisfies $\mathbb{E}_{\xi}\xi^{*}_{i}\xi_{j}=\delta_{ij}$, including Gaussian distribution, Rademacher distribution, etc [70]. In practice, we compute $\sum_{l}\left[\left(M^{\dagger}M\right)^{-1}\right]_{il}\left[M^{\dagger}\xi\right]_{l}\xi^{*}_{j}$, instead, where $R:=M^{\dagger}\xi$ is numerically cheap to compute and $\left[M^{\dagger}M\right]^{-1}R$ can be computed generically using a pCG algorithm. In this way, the Green’s function can be obtained, and other more complicated correlation functions can be further obtained from Green’s function via Wick decomposition, as has been practiced in many other previous DQMC simulations [71, 72].
Note that in practice, when numerically computing the Green’s functions, we utilize the translational invariance and take the average over space and imaginary time, i.e. $C(i,j)=C(0,\Delta)$. Here, $i,j$ are site indices of the cubic spatial-temporal lattice, $\Delta:=j-i$, and $C(i,j)$ is either two-point, four-point or eight-point Green’s functions. Apparently, naively taking the average
$C(\Delta)=\frac{1}{N}\sum_{i}C(i,i+\Delta)$ (14)
has a time complexity of $O(N^{2})$, where $N=V_{s}N_{\tau}$. Here, we follow the Ref. [70], and apply the fast Fourier transformation (FFT) instead of directly eevaluating this average.
As introduced above, for a given bosonic field configuration, and a random vector $\xi$, an unbiased estimation of the two-point Green’s function is $\hat{G}_{ij}=\left[G\xi\right]_{i}\xi^{*}_{j}$. Thus, an unbiased estimation of any correlation function $C(i,j)$, either the four-point Green’s function or the eight-point Green’s function, can be uniformly written in the form of
$\hat{C}(i,j)=a_{i}b_{j},$ (15)
where $a$ and $b$ are two vectors. For example, if $C(i,j)=G_{i,j}G_{i,j}$, then for an unbiased estimation of $\hat{C}(i,j)$, $a_{i}=(G\xi)_{i}(G\xi^{\prime})_{i}$ and $b_{j}={\xi^{*}_{j}}{\xi^{*}_{j}}^{\prime}$. Then, the average of $\hat{C}(i,i+\Delta)$ takes the convolution form, and FFT is readily applied,
$\hat{C}(\Delta)=\frac{1}{N}\sum_{i}a_{i}b_{i+\Delta}=\mathcal{F}^{-1}[\mathcal{F}[a]_{-k}\cdot\mathcal{F}[b]_{k}](\Delta),$ (16)
where $\mathcal{F}$ is FFT operation. Then the time complexity is reduced to $O(N\log N)$, with $N=V_{s}N_{\tau}$. Moreover, as introduced in Ref. [70], each random vector among $N_{\textmd{rv}}$ random vectors can serve as either $\xi$ or $\xi^{\prime}$ in the example above. Thus, totally $f(N_{\textmd{rv}})=\binom{N_{\textmd{rv}}}{2}$ random vectors can be effectively used in the stochastic estimation. In our benchmark, $N_{\textmd{rv}}=40$ is sufficient to reduce the error of fermionic bilinear’s correlation function to a moderate level.
II.4 CG method and preconditioner
As introduced above, the key step in reducing the cubic complexity of computing the determinant $\det M$ or the inversion $M^{-1}$ is to convert it to the evaluation of $\left[M^{\dagger}M\right]^{-1}v$, which is equivalent to solving a linear system $OX=v$. Here, $O=M^{\dagger}M$ is a Hermitian positive semi-definite matrix; thus, the equation can be efficiently solved by the iterative preconditioned conjugate gradient (pCG) method [53].
The pCG iteration can be summarized as follows. Given the linear equation $OX=v$ and the preconditioner $\tilde{O}^{-1}$, initialize the vectors $X_{0}=\mathbf{0}$, $r_{0}=v-OX_{0}$, $d_{0}=r_{0}$ and the scalar $\zeta_{0}=\left\langle r_{0},\tilde{O}^{-1}r_{0}\right\rangle/\left\langle d_{0},Od_{0}\right\rangle$. Then, for $n=0,1,2,\ldots$ until convergence,
$X_{n+1}$ $=X_{n}+\zeta_{n}d_{n}$
$r_{n+1}$ $=r_{n}-\zeta_{n}Od_{n}$
$\rho_{n}$ $=\left\langle r_{n+1},\tilde{O}^{-1}r_{n+1}\right\rangle/\left\langle r_{n},\tilde{O}^{-1}r_{n}\right\rangle$ (17)
$d_{n+1}$ $=\tilde{O}^{-1}r_{n+1}+\rho_{n}d_{n}$
$\zeta_{n+1}$ $=\left\langle r_{n+1},\tilde{O}^{-1}r_{n+1}\right\rangle/\left\langle d_{n+1},Od_{n+1}\right\rangle,$
where $\langle\cdot,\cdot\rangle$ is the inner product. The time complexity of this algorithm is decided by the matrix-vector multiplication time complexity in $\tilde{O}^{-1}r$ and $Od$, and the number of pCG iterations.
In large-scale computations of pCG algorithm, the application of preconditioner is crucial in reducing the number of iterations required to reach convergence. However, the choice of preconditioner is usually problem-specific; there is not a universal way of obtaining a proper preconditioner for any linear system [53]. In principle, a good choice of preconditioner needs to balance the accuracy and efficiency. The accuracy is to measure how well the preconditioner approximates the inverse of the $O$ matrix in $OX=v$, i.e., the preconditioner $\tilde{O}^{-1}\approx O^{-1}$. In the most ideal case, if $\tilde{O}^{-1}=O^{-1}$, then it would require only one iteration to reach convergence and obtain $X=O^{-1}v$. But if the preconditioner were too complex to obtain or to utilize in pCG iterations, then the pCG process would not be accelerated.
In our case, the preconditioner $\tilde{O}^{-1}$ is obtained in the following procedure. First, for a given matrix $O=M^{\dagger}(\phi)M(\phi)$ for some generic fixed bosonic field $\phi$, we apply the incomplete Cholesky decomposition to obtain a lower triangular matrix $L(\phi)$, such that $O\approx L^{\dagger}(\phi)L(\phi)$. Then, we apply the Neumann series expansion to approximate the inverse of $L$ as
$L^{-1}$ $\approx 1+(1-L)+(1-L)^{2}+\ldots,$ (18)
for $20$ iterations. Finally, we truncate the entries with values smaller than $1\times 10^{-3}$, and obtain the preconditioner $\tilde{O}^{-1}=L^{-1}(\phi)L^{-\dagger}(\phi)$. Here, all computations, including incomplete Cholesky decomposition and triangular matrix multiplications, have time complexity $O((V_{s}N_{\tau})^{2})$, which is numerically affordable. Most importantly, we find that the preconditioner in our problem is endowed with good generalizibility, in the sense that the preconditioner obtained at a generic bosonic field $\phi$ is applicable to accelerate the pCG iteration at any other bosonic fields throughout the Monte Carlo simulation. So in practice, we just use $\pi$-flux bosonic configuration to compute the preconditioner once and apply it globally throughout the MC simulation.
In essence, such generalizability in the preconditioner is because in our problem, as shown in Eq. (5), the $M$ matrix is very weakly modulated by the bosonic field, as long as $\Delta_{\tau}$ is small enough. Our computation shows that as long as $\Delta_{\tau}\leq 0.1$, the pre-computed preconditioner has good generalizability. But if $\Delta_{\tau}\gtrsim 0.12$, then the pre-computed global preconditioner will fail and have a negative impact on the convergence of pCG iterations.
[…]
II.5 Customized CUDA kernel for matrix-vector multiplication
The mathematical operations with the highest cost in the pCG iterations are the two matrix-vector multiplications, namely $M^{\dagger}Mv$ and $\tilde{O}^{-1}v$ multiplication. It turns out that these two operations can benefit significantly from writing customized CUDA kernels, which effectively exploit the specific sparsity pattern of the matrices.
II.5.1 Matrix-free representation of $M^{\dagger}Mv$ multiplication
A naive implementation of this $M^{\dagger}Mv$ operation would first compute $M^{\dagger}M$, and store it as a sparse matrix, and then time it with the vector $v$. Generally speaking, the multiplication of a sparse matrix $M^{\dagger}M$ with a vector $v$ has the time complexity of $O(nnz)$ where nnznnz is the number of non-zero elements in the matrix $M^{\dagger}M$, which can reach the scale of $O(N^{2})$ when the matrix is not sparse enough. However, instead of the sparse-matrix representation, the definition of the $M$ matrix in Eq. (5) admits a matrix-free representation in the matrix-vector multiplication, which is more efficient [73]. This technique helps achieve a 2.9 – 3.8 times speedup on top of the baseline implementation, as will be shown next.
In the definition of the $M$ matrix in Eq. (5), we have $B_{\tau}=e^{V_{\tau}}$, with $\left[V_{\tau}\right]_{ij}=\Delta_{\tau}e^{i\phi_{ij,\tau}}$. Then, due to the geometry of the square lattice, one can symmetrically decompose the $V_{\tau}$ matrix into four families, as denoted as the four bond colors in Fig. 1:
$B\approx e^{V_{4}/2}e^{V_{3}/2}e^{V_{2}/2}e^{V_{1}/2}e^{V_{1}/2}e^{V_{2}/2}e^{V_{3}/2}e^{V_{4}/2}$ (19)
($\tau$ indices are omitted), which is also known as the symmetrized checkerboard decomposition [6]. This approximation amounts to ignoring the non-commutativity between different families of $V$ matrices, and thus has the same error contribution as the Trotter error. But within a bond family, all hopping terms commute with each other. Therefore, the matrix $V_{\textmd{fam}_{n}}$ can be written as a direct sum of $2\times 2$ matrices:
$e^{V_{\textmd{fam}_{n}}}=e^{\Delta\tau\bigoplus_{ij\in\textmd{fam}_{n}}T_{ij}},$ (20)
where $T_{ij}=\left[\begin{array}[]{cc}0&e^{i\phi_{ij}}\\ e^{-i\phi_{ij}}&0\end{array}\right],$ which has property of $T_{ij}^{2}=\mathbf{1}$. Then within a bond family $\textmd{fam}_{n}$, the exponentiation can be readily evaluated:
$e^{V_{\textmd{fam}_{n}}}$ $=\bigoplus_{ij\in\textmd{fam}_{n}}e^{\Delta_{\tau}T_{ij}}$ (21)
$=\bigoplus_{ij\in\textmd{fam}_{n}}\left(\mathbf{1}\cosh\Delta_{\tau}+T_{ij}\sinh\Delta_{\tau}\right)$ (22)
$=\left[\begin{array}[]{cccc}\ddots&&&\\ &\cosh\Delta_{\tau}&e^{i\phi_{ij}}\sinh\Delta_{\tau}&\\ &e^{-i\phi_{ij}}\sinh\Delta_{\tau}&\cosh\Delta_{\tau}&\\ &&&\ddots\end{array}\right]_{\textmd{fam}_{n}}.$ (27)
It is then clear that the $Bv$ multiplication can be written as a composite of serial sparse-matrix-vector multiplication $e^{V_{\textmd{fam}_{n}}}v$, which is straightforwardly parallelizable in GPU, and has a time complexity of $O(V_{s}N_{\tau})$. Furthermore, in the series of $u=e^{V_{\textmd{fam}_{n}}}v$ matrix-vector multiplication, the intermediate input vector $v$ and output vector $u$ can be put into the shared memory of GPU, which spares the communication cost between the high-bandwidth memory (HBM) and the on-chip static random-access memory (SRAM) [74].
II.5.2 Preconditioner-vector multiplication $\tilde{O}^{-1}v$
Figure 3: Banded preconditioner matrix and its CUDA implementation. (a) The plot shows the banded sparsity pattern of the preconditioner $\tilde{O}^{-1}$, where the nonzero elements of the first 500 by 500 entries are displayed. The bandwidth is 6. This preconditioner is obtained for a lattice of $N_{\tau}\times L\times L=60\times 6\times 6$. Two neighboring diagonal lines are separated by a distance of $V_{s}=L\times L$ which is $36$ in this illustration. (b) Schematic of the matrix-free preconditioner-vector multiplication $u=\tilde{O}^{-1}v$, which is illustrated with a banded preconditioner matrix with bandwidth $1$. A block of threads are assigned to process the rows of the preconditioner matrix sandwiched by the orange dashed line, with one thread processing one row. The chunk of elements in $v$ involved in this computation are denoted by blue dots, where light blue ones are called the halo elements. In our CUDA kernel implementation, the chunk of $v$ represented by the blue dots are copied into shared memory, which utilizes the local caching of the GPU and reduces the communication overhead.
As shown in the Fig. 3 (a), our preconditioner obtained in the procedure introduced in Sec. II.4 has a fixed banded sparsity pattern in the $\tau$ dimension, with a bandwidth $6$. Here we will show how to exploit such sparsity pattern to accelerate preconditioner-vector multiplication $\tilde{O}^{-1}v$.
A standard way to write a CUDA kernel for such banded sparse matrix-vector multiplication is illustrated in Fig. 3 (b), exemplified here with a bandwidth of 1 [73, 55]. In this illustration, a block of threads is assigned to process the rows of the matrix, with one thread corresponding to one row of the matrix. In turn, in the matrix-vector multiplication of these rows requires an input of a chunk of the vector $v$ shown as the blue dots in the figure. Apparently, the thread block size is equal to the number of rows being processed and is equal to the vector chunk size minus the number of the so-called halo elements (the light blue dots in Fig. 3 (b)). Thus, the computation is well parallelized. Besides, before the computation, the chunk of the vector is first copied to the shared memory. In this way, the neighboring elements in the vector can be simultaneously read by the neighboring threads. Such local caching has thus been exploited to effectively minimize the communication overhead between HBM and the on-chip SRAM, similar to the $M^{\dagger}Mv$ kernel above. The time complexity of this preconditioner-vector multiplication $\tilde{O}^{-1}v$ is linear with respect to $N_{\tau}$ and $V_{s}$.
II.5.3 Speedup of customized CUDA kernel
Figure 4: Speedup of our customized CUDA kernel. The speedup achieved by customized CUDA kernel for different system sizes $L$, with $V_{s}=L^{2}$ and $N_{\tau}=10L$. The blue axis on the left represents the latency of the Monte Carlo sweep which is average time cost per generated sample. The light blue line represents the baseline performance, where the PyTorch sparse-matrix-vector multiplication is applied to compute $M^{\dagger}Mv$ and $\tilde{O}^{-1}v$. The dark blue line is the latency of the customized CUDA kernel, where the computations $M^{\dagger}Mv$ and $\tilde{O}^{-1}v$ are replaced with optimized CUDA kernel. The red dashed line is the latency ratio, which shows the speedup between the two implementations for each system size.
Fig. 4 shows the speedup of the above two customized CUDA kernels over the naive implementations. The baseline for implementing $M^{\dagger}Mv$ and $\tilde{O}^{-1}v$ is to directly call PyTorch sparse-matrix-vector multiplication function, where $M^{\dagger}M$ and $\tilde{O}^{-1}$ are stored as sparse compressed row storage matrices. The hardware is NVIDIA l40s GPU. Our test shows that the customized CUDA kernel achieves up to 2.9 – 3.8 times speedup across all system sizes.
II.6 CUDA graph acceleration
Figure 5: Illustration of CUDA Graph acceleration. (a) The upper half illustrates the naive GPU implementation without using CUDA Graph. The blue boxes are a series of operations (kernels) to perform in GPU, which need to be launched one by one in CPU, represented by the orange rectangles. (b) The lower half shows the mechanism of CUDA Graph acceleration, where the series of operations (kernels) are first compiled together, which happens in the Build Graph phase, and then launched only once in the execution phase. By comparing the wall-clock time between the upper and the lower half, one can see the achieved speedup.
Figure 6: The speedup achieved by CUDA Graph. For different system sizes $L$, with $V_{s}=L^{2}$ and $N_{\tau}=10L$, the blue axis on the left represents the latency of the Monte Carlo sweep which is average time cost per generated sample. The light blue line represents the baseline performance with CUDA Graph turned off, while the dark blue line is the latency when CUDA Graph is turned on. The red dashed line is the latency ratio, which shows the speedup between the two implementations for each system size.
CUDA graph is another important technique that helps achieve significant speedup in the numerical simulation, which is around 2-5 times on moderate lattice sizes (around $L<30$) and 1-2 times on large system sizes, compared to the baseline implementation. This technique was not introduced until very recently, and has been widely applied in a variety of large-scale GPU-accelerated computations and systems ever since [75, 76, 77, 78, 79, 80, 81].
The principle of the CUDA graph is illustrated in Fig. 5. Without using CUDA graph, if we have a stream of GPU operations to process, generically the operations, implemented as CUDA kernels, will need to be launched one by one in the CPU and then executed in the GPU. As illustrated in Fig. 5, such repetitive launch of CUDA kernels in CPU leads to significant total launching overhead, and also the intermittent execution of kernels causes the GPU to idle between two executions; the total wasted time is comparable to the GPU effective utilization time. But if CUDA graph is applied, what it does differently is to first launch and execute a stream of kernels in a warm-up phase, and record (compile) the computation into a computation graph. Then, in the actual execution phase, the whole computation graph is launched only once and all operations are executed altogether in the GPU. Thus, the speed-up is achieved by reducing the total kernel launching overhead.
However, this technique has the following major limitations that need to be considered in application. The first is that the storage of the entire computation graph increases the GPU memory usage. So, GPU memory capacity can be a bottleneck when scaling up the computation. The second limitation is that the recorded CUDA graph is static, in the sense that once the CUDA graph is recorded, all the subsequent executions of the graph cannot early exit and that the input array must remain the same size and shape at the same memory address, while only its entry values can vary in each CUDA graph launch.
An important consequence of this static graph limitation is that, if the whole pCG iterations are recorded as a CUDA graph, then the number of iterations will be fixed, which cannot be early terminated when the error reaches a fixed precision level. So in our benchmark, instead of fixing the precision level and testing the convergence iteration number, we fix the number of iterations and test the precision.
Despite the limitation, the speedup of this technique is important. Fig. 6 shows the latency of a HQMC sweep for various lattice sizes and compares two scenarios between CUDA Graph on and off. In our implementation, we convert the whole fermionic force computation, as well as the stochastic estimation of fermionic Green’s function, into CUDA Graph. Fig. 6 shows that the execution is around 2-5 times faster for $L<30$ and 1-2 times faster for larger lattice sizes. The reason that the speedup from CUDA Graph decreases with increased system size, is because as the system size increases, the GPU kernel execution time is increasingly intense. Thus, the kernel launching time becomes relatively small compared to the kernel execution time. The speed up from CUDA Graph acceleration is mainly from the saving of kernel launching latency, so as system size increases, the speedup becomes smaller.
The speedup from CUDA Graph is orthogonal to the speedup obtained from the customized CUDA kernels in Sec. II.5, thus the speedups are multiplicative.
II.7 Complexity analysis
Figure 7: The $\bm{O(N_{\tau}V_{s})}$ computational complexity of HQMC. The scaling behavior of the HQMC latency and stochastic estimation latency as functions of $N_{\tau}V_{s}\propto L^{3}$, where we have set $N_{\tau}=\beta/\Delta_{\tau}=10L$ throughout the test. The latency is defined as the time spent to generate a gauge sample for HQMC, or the time to compute a fermionic observable from a gauge sample for SE (the SE latency here is tested on spin-spin correlation using $N_{\textmd{rv}}=40$). The blue and orange line are the linear fitting of five large-size points. They show that the HQMC algorithm has a time complexity of $O((L^{3})^{0.9})$, while the stochastic estimation has a time complexity of $O((L^{3})^{1.3})$. As a contrast, the DQMC has a time complexity of $O(N_{\tau}V_{s}^{3})=O(L^{7})$, which is plotted as the green dashed line. The starting point of the green dashed line is $0.38$ second per sweep at $L^{3}=10^{3}$, obtained from DQMC benchmark.
Now we give an overview of the complexity analysis of our HQMC. As introduced above, the main iteration of generating a sample in HQMC consists of first $N_{\textmd{leapfrog}}$ steps of Leapfrog integration as shown in Eq. (13). Within each Leapfrog step, in computing the fermionic force, $N_{\textmd{pCG}}$ steps of pCG iteration shown in Eq. (17) are required to converge to a given precision. Finally, within each pCG iteration, two matrix-vector multiplications are involved, which have time complexity of $O(N_{\tau}V_{s})$. So the overall time complexity of generating a sample is
$T_{\textmd{sample}}=N_{\textmd{leapfrog}}\cdot N_{\textmd{pCG}}\cdot O(N_{\tau}V_{s}).$ (28)
As introduced above, $N_{\textmd{leapfrog}}$ and $N_{\textmd{pCG}}$ have constant time complexity. So the time complexity of generating a sample is $T_{\textmd{sample}}=O(N_{\tau}V_{s})$, which is shown as the blue line in Fig. 7.
The stochastic estimation for a given sample has similar scaling of time complexity. For a given random vector, a pCG iteration is also called in the stochastic estimation process. So the time complexity is $O(f(N_{\textmd{rv}})N_{\tau}V_{s})$, where $N_{\textmd{rv}}$ is the number of random vectors, and $f(N_{\textmd{rv}})$ is a polynomial function of $N_{\textmd{rv}}$, which is fixed to be $\binom{N_{\textmd{rv}}}{2}$ in this benchmark of fermionic bilinears’ correlation functions. Such scaling behavior with respect to $N_{\tau}V_{s}$ is shown as the orange line in Fig. 7.
As shown in Fig. 7, both the sample generation and stochastic estimation display nearly linear scaling behavior with respect to $N_{\tau}V_{s}$ as expected, so does the overall time complexity. Moreover, the stochastic estimation has slightly larger scaling power $0.905$ than sample generation $1.313$, and also larger latency at large system sizes. This is because stochastic estimation needs to simultaneously process $f(N_{\textmd{rv}})$ input vectors, while sample generation only needs to process one. Thus, the stochastic estimation has much stronger arithmetic intensity, and the latency scales faster with respect to system size. Nevertheless, Fig. 7 also shows a green dashed guideline which is the computational complexity of DQMC. At system size of $L^{3}\sim 10^{5}$, the two algorithms already have two orders of magnitude latency difference, which shows the advantage of the HQMC.
III Results
Here we show the results from our GPU-accelerated HQMC simulations, we focuse on the non-compact gauge field case, and will present results on the compact case in SM [64]. For small system sizes, we benchmark the HQMC with the DQMC and find consistency at the quantitative level, up to sizes of $160\times 16\times 16$ ($N_{\tau}\times L\times L$). We then use HQMC to push for the thermodynamic limit, up to $660\times 66\times 66$, which is more than $50$ times larger in the space-time volume limit of DQMC, and the data clearly reveal the power-law correlations in different fermion bilinear operators and the flux operator inside the $U(1)$ DSL phase.
III.1 Physical observables
We first introduce the fermionic observables used to characterize and detect the $U(1)$ DSL phase. In our problem of QED3 $U(1)$ gauge-field coupled fermions, one caution we need to take is the gauge invariance of an observable. For example, the single-particle Green’s function is not gauge-invariant, and to make it gauge-invariant, one will require an extra integral of the $U(1)$ phase along the path connecting the two fermions. On the other hand, the fermion bilinear operators are gauge-invariant, and we measure the following two correlation functions of fermion bilinears – spin-spin and bond-bond correlation functions – as
$C_{\textmd{S}}^{\alpha\beta}(i,j)$ $=\langle S^{\alpha\beta}(i)S^{\beta\alpha}(j)\rangle$ (29)
$C_{\textmd{B}}(i,j)$ $=\langle B_{i}B_{j}\rangle-\langle B_{i}\rangle\langle B_{j}\rangle,$ (30)
where the spin operator is defined as $S^{\alpha\beta}(i)=\psi_{i\alpha}^{\dagger}\psi_{i\beta}-\frac{1}{N_{f}}\delta_{\alpha\beta}\sum_{\gamma}\psi_{i\gamma}^{\dagger}\psi_{i\gamma}$ , bond operator along the nearest-neighbor bond in $\hat{x}$ direction is defined as $B_{i}=\sum_{\alpha}(\psi^{\dagger}_{i,\alpha}e^{i\phi_{i,i+\hat{x}}}\psi_{i+\hat{x},\alpha}+\text{h.c.})$, and $\alpha,\beta\in\left\{\uparrow,\downarrow\right\}$ are spin flavor indices. Here, the spin operators are written as antisymmetric self-conjugate representation [1], and are related to the normal spin operators definition as $S^{\uparrow\downarrow}=S^{+},S^{\downarrow\uparrow}=S^{-},S^{\uparrow\uparrow}=S^{z},S^{\downarrow\downarrow}=-S^{z}$. The corresponding structure factors are defined as the Fourier transformations of the correlation functions $\chi_{\textmd{S, B}}(\mathbf{k})=\frac{1}{V_{s}^{2}}\sum_{ij}C_{\textmd{S, B}}(i,j)e^{i\mathbf{k}\cdot(\mathbf{r}_{i}-\mathbf{r}_{j})}$. The spin form factor $\chi_{S}(\mathbf{k})$ evaluated at $X=(\pi,\pi)$ point will be used to characterize the antiferromagnet (AFM) order, while the bond structure factor $\chi_{B}(\mathbf{k})$ evaluated at $M=(\pi,0)$ point, will be used to characterize the columnar VBS order.
Other than the gauge invariant fermionic observables, we also measure the correlation of the magnetic flux operator $\theta_{\square}(\tau)$, with a focus on its imaginary time dynamics. Since, in the static limit ($J=0$) of the QED3, the ground energy magnetic configuration is known to be $\pi$-flux on a square lattice [61]. When $J$ is turned on, the fluctuation of magnetic flux is centered also around $\pi$-flux. So, to study the correlation of the magnetic fluxes, we measure the correlation of its sine value instead, in order to have a clean background, i.e. $\langle\sin\theta_{\square}\rangle=0$. We define
$C_{\textmd{flux}}(\tau)$ $=\frac{1}{V_{s}}\sum_{\square}\langle\sin\theta_{\square}(\tau)\sin\theta_{\square}(0)\rangle.$ (31)
III.2 The operator scaling dimensions in non-compact QED3
As mentioned in the Introduction (Sec. I.1), the DSL with non-compact QED3 has been analyzed extensively and is believed to describe a scale-invariant conformal field theory (CFT). Consequently, power law behavior is expected for all gauge-invariant observables. At our setting of $N_{f}=2$, the CFT is expected to have emergent $SU(4)$ symmetry (with two cones in the BZ), one expects the scaling dimension of the fermion bilinears or mass terms: a singlet $\Delta_{s}\approx 2.3$ and a set of adjoints $\Delta_{\text{adj}}\in(1.4,1.7)$, where there are 15 of them ranging over the $SU(4)$ generators. These scaling dimensions are obtained from the $\epsilon$-expansion from $d=4-2\epsilon$ to 3 [18]. We note that, the spin operator and the bond operator mentioned above in Eqs. (29) and (30) are among the $15$ adjoint fermion bilinear terms and consequently they are expected to share the same power-law ($\sim 1/r^{2\Delta_{\text{adj}}}$) when we compute their correlation functions [3, 6] and it is the indeed the case as we show in the next section.
Moreover, in the gauge sector, one expects that the $U(1)$ DSL should exhibit flux-flux correlations dictated by $\Delta_{J}=2$, which implies the long-time correlations of the local flux should behave as $\langle\sin\theta_{\Box}(\tau)\sin\theta_{\Box}(0)\rangle\sim\tau^{-4}$. In contrast, the flux correlation in a non-compact $U(1)$ gauge field theory has a power-law dynamic correlation of $\langle\sin\theta_{\Box}(\tau)\sin\theta_{\Box}(0)\rangle\sim\tau^{-3}$, which is non-conformal but instead from a Goldstone mode of spontaneously broken magnetic $U(1)$ symmetry, related to flux conservation [82]. From our HQMC simulation, the $\tau^{-4}$ behavior of flux correlation is seen at large space-time volume.
Figure 8: The spin correlation function for non-compact QED3. The log-log plot of the spin-spin correlation $C_{S}^{\uparrow\downarrow}(r,0)$ as a function of distance $r$ for non-compact QED3in Eqs. (1) and (3) with $K=0$ and $J=1.25$ , for various cubic lattice sizes shown as $N_{\tau}\times L^{2}$ in the legend. Notice that $C_{S}^{\uparrow\downarrow}=C_{S}^{\downarrow\uparrow}$, so we only need to measure $C_{S}^{\uparrow\downarrow}$ here. To avoid the even-odd oscillation in the finite-size data, we only plot the data of the distance $r=\textmd{odd}$ points. The gray square, triangular and diamond dots show the results computed from the DQMC to benchmark the HQMC result of the corresponding lattice sizes. At the large distance and the large system sizes, the data are manually fitted by the black solid line, which displays a power law decay as $1/r^{3.0}$.
Figure 9: The bond correlation function for non-compact QED3. The log-log plot of the bond-bond correlation (along $\hat{x}$ direction) as a function of distance $r$ for non-compact QED3 in Eqs. (1) and (3) with $K=0$ and $J=1.25$ , for various lattice sizes shown as $N_{\tau}\times L^{2}$ in the legend. At the large distance and the large system sizes, the data are manually fitted by the black line, which displays a power law decay as $1/r^{3.3}$.
Figure 10: The flux correlation function for non-compact QED3. The log-log plot of the flux-flux correlation as a function of imaginary time $\tau$ for non-compact QED3 with $K=0$ and $J=1.25$ Eqs. (1) and (3), for various lattice sizes, shown as $N_{\tau}\times L^{2}$ in the legend. At the large distance and the large system sizes, the data are manually fitted by the black line, which displays a power law decay as $\tau^{-4}$. The black dashed line shows the $\tau^{-3}$ power law behavior for comparison.
III.3 Numerical results
In this section, we analyze the numerical results of the fermionic correlation functions and the gauge flux correlation functions in non-compact QED3 in Eqs. (1) and (3) as introduced above, with a focus on their scaling dimensions and how they shed light on the existence of the $U(1)$ DSL phase.
Fig. 8 shows the log plot of the spin-spin correlation $C_{S}^{\uparrow\downarrow}(r,0)$ as a function of real space distance $r$ at $J=1.25,K=0$, for various lattice sizes. As the system size increases from $L=12$ all the way to $L=60$, the spin correlation curves monotonically unroll, and approach a straight envelope line at large distances. This indicates a power law decay behavior. By manually fitting this envelop curve shown as the dark solid line, we obtain the exponent of the power law decay to be around $3.0$. This agrees with the scaling dimension of the fermion bilinear operators $2\Delta_{\textmd{adj}}\in(2.8,3.4)$ predicted by the CFT, where the fermion bilinear operators serve as the adjoint representation of the emergent $SU(4)$ symmetry. We note that around $10^{-5}$ in Fig. 8, the data fall below the noise level. So the manual fitting is performed mainly with the few large-distance data $\gtrsim 10^{-5}$ at large system sizes. In Fig. 8, we also plot the DQMC results for several small lattice sizes to benchmark the HQMC results, and they well agree with each other. These results are complemented by the benchmark data in Sec.II of the Supplemental Materials (SM) [64].
Fig. 9 shows log plot of the bond-bond correlation $C_{B}(r,0)$ as a function of real space distance $r$ at $J=1.25,K=0$, for various lattice sizes. The data of the bond correlation generally have a larger variance than the spin correlation function shown in Fig. 8, which is due to the explicit gauge field dependence in the bond operator $B_{i}$. But such variance from gauge field fluctuation seems to be alleviated at large system sizes, and the noise level decreases to around $5\times 10^{-5}$ as the system size increases to $N_{\tau}\times L^{2}=600\times 60^{2}$. Furthermore, for system sizes $L\gtrsim 40$, the decaying behavior of the bond-bond correlation is stable around a straight line, which again indicates a power law decay. The manual linear fitting shown as the dark solid line yields a power of around $3.3$, which again agrees with the scaling dimension of the fermion bilinear operators $2\Delta_{\textmd{adj}}\in(2.8,3.4)$.
As analyzed in the above section and shown in Fig. 8 and Fig. 9, the fact that both spin and bond correlation functions display power law decay with powers consistent with the scaling dimension of the CFT fermion bilinear operators supports the existence of the $U(1)$ DSL phase in non-compact QED3 at $J=1.25$. This also agrees with a recent DQMC study on the same model [82].
Fig. 10 shows the log plot of the magnetic flux correlation function $C_{\textmd{flux}}(\tau)$ as a function of imaginary time $\tau$ for various lattice sizes, at $J=1.25,K=0$. It can be seen that, as system size increases, the curves become increasingly linear before entering the noise level at around $2\times 10^{-3}$. Such linear tendency is manually fitted with the dark solid line, which yields a power of around $4.0$. In contrast, the black dashed line shows the $\tau^{-3}$ behavior, which originates from the non-conformal flux-flux correlation in a $U(1)$ gauge field theory [83]. This is clearly distinct from the $\tau^{-4}$ behavior originating from the conserved current’s scaling dimension $\Delta_{J}=2$, as mentioned above. Thus, the power law decay in the flux correlation function with power $4.0$ again indicates the existence of $U(1)$ DSL phase in $J=1.25$ in the non-compact QED3.
IV Discussions
In this work, we developed a scalable GPU-accelerated hybrid quantum Monte Carlo (HQMC) algorithm for simulating $U(1)$ gauge field coupled to fermions in (2+1)d. The technical advances achieved here are threefold:
  1. 1.
    the construction of a problem-specific preconditioner that enables rapid convergence of the conjugate gradient solver;
  2. 2.
    the design of customized CUDA kernels that exploit the sparsity and structure of the matrix-vector operations intrinsic to the model;
  3. 3.
    the integration of CUDA graph acceleration to minimize kernel launch overhead and maximize GPU utilization.
These optimizations collectively allow us to reach a linear scaling in computational complexity with respect to space-time volume, i.e. $O(N_{\tau}V_{s})$, and to simulate unprecedentedly large system sizes with ease.
Beyond the computational advances, our large-scale simulations have enabled us to investigate the non-compact QED3 with better scrutiny. By systematically analyzing the spin-spin and bond-bond correlation functions, we extract the scaling dimensions of the corresponding fermion bilinear operators. Our results demonstrate that both observables exhibit power-law decay with exponents consistent with scaling dimension of the adjoint representation bilinear terms of the emergent SU(4) symmetry computed from field-theoretical approach. Furthermore, the flux-flux correlation function reveals a $\tau^{-4}$ decay at long times, reflecting the scaling dimension $\Delta_{J}=2$ of the conserved current. These results provide direct evidence for the conformal nature of the $U(1)$ Dirac spin liquid phase in the non-compact QED3 model.
Since the HQMC gives the full space-time correlation of the fermion bilinear operators, one immediate future direction is to investigate the spectral information inside the $U(1)$ DSL phase from the stochastic analytic continuation of the HQMC data [82]. As we have both large $N_{\tau}$ and $V_{s}$, the momentum and energy resolution of the obtained spectra (for example the magnetic spectra from spin-spin correlation function) will be much higher than those from the DQMC [82]. In this way, the weight distribution and the momentum dependence of the spinon continuum spectra of $U(1)$ DSL can be revealed, and hopefully connection with experiment results on triangular lattice antiferromagnets YbZn2GaO5 [35, 36] and the $A$-YbSe2 delafossites [37, 38] and kagomé antiferromagnet YCu3(OD)6Br2[Br1-x(OD)x[39, 40, 41, 42, 43, 44, 45] can be made.
Moreover, by adding magnetic field to the QED3 model, one can investigate how the flavor chemical potential will change the $U(1)$ DSL phase and understand the emergent gauge flux in magnetized Dirac spin liquids. For example, recent DQMC work [82] and analytic works [84, 85] have explored the various symmetry-breaking phases out of $U(1)$ DSL under magnetic field. In particular, a novel chiral flux phase, where the system spontaneously generates a net emergent orbital magnetic flux (deviated from the $\pi$-flux at zero field case) has been seen from smaller size DQMC simulation. This spontaneous orbital magnetic flux changes the behavior of the spinons, and make them form the relativistic Landau levels — energy levels normally seen in electrons moving in a magnetic field. Remarkably, the system also exhibits a gapless photon-like mode, a signature of the coherent emergent gauge field, which can be potentially detected in inelastic neutron scattering experiments [82]. Our GPU-accelerated HQMC algorithm can provide systematic results of such interesting phase and rich diagram, as a function of magnetic field, gauge field fluctuations, as well as the temperature axes. One possible direction is to investigate the effect of different filling of the spinon Landau levels in the magnetized DSL phase [86], and establish possible connections to experimentally observed unusual magnetic oscillations in the YCu3(OD)6Br2[Br1-x(OD)x] kagomé antiferromagnet [45].
Taken together, our work establishes a robust and efficient computational framework for exploring the physics of strongly coupled quantum field theories at large scales, and provides new quantitative insights into scaling dimensions of the operators in the $U(1)$ Dirac spin liquid. This technique opens an avenue to future investigations of phase transitions and symmetry-breaking phenomena in related models, and hopefully, making closer connection with the active on-going experimental efforts in finding the $U(1)$ DSL state in quantum magnets and the theoretical efforst in understanding the strongly coupled conformal field theory in 2+1 dimensions in high-energy physics.
Acknowledgements.
We thank U.F.P. Seifert, O.A. Starykh and L. Balents for the stimulating discussion on the physics of $U(1)$ DSL and QED3, and the collaboration on a related work [82]. We thank M. Ulybyshev and F. Assaad for inspiring communications on hybrid Monte Carlo [87]. KXF thanks Natalia Perkins for the collaboration in a related project. KXF, CC and ZYM acknowledge the support from the Research Grants Council (RGC) of Hong Kong (Project Nos. AoE/P-701/20, 17309822, HKU C7037-22GF, 17302223, 17301924), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A HKU703/22). We thank HPC2021 system under the Information Technology Services and the Blackbody HPC system at the Department of Physics, University of Hong Kong, as well as the Beijng PARATERA Tech CO.,Ltd [88] for providing HPC resources that have contributed to the research results reported within this paper.
Supplemental Material
Scalable Hybrid quantum Monte Carlo simulation of QED3 coupled to fermions on GPU
In this supplemental material, we provide the technical details concerning the QMC measurements of conserved currents (Sec. SII) and the finite size scaling of the order parameters at the easy-plane DQCP (Sec. SIII).
SI. Calculation of Fermionic Force for QED3
The fermionic force in the Hamiltonian dynamics can be derived in the following way,
$F_{i,\tau}=$ $\frac{\partial}{\partial\phi_{i,\tau}}\eta^{\dagger}\left(M^{\dagger}(\phi)M(\phi)\right)^{-1}\eta$ (S1)
$=$ $-\eta^{\dagger}\left(M^{\dagger}(\phi)M(\phi)\right)^{-1}\left[M^{\dagger}(\phi)\frac{\partial}{\partial\phi_{i,\tau}}M(\phi)+\left(\frac{\partial}{\partial\phi_{i,\tau}}M^{\dagger}(\phi)\right)M(\phi)\right]\left(M^{\dagger}(\phi)M(\phi)\right)^{-1}\eta,$ (S2)
where $i$ is spacial lattice site position. Here, the $\left[M^{\dagger}M\right]^{-1}\eta$ can be computed by the pCG algorithm as introduced in the main text, with complexity of $O(N_{\tau}V_{s})$. This can be expressed as a vector $X=\left[X_{1}^{\intercal},X_{2}^{\intercal},\ldots,X_{N_{\tau}}^{\intercal}\right]^{\intercal}:=\left[M^{\dagger}M\right]^{-1}\eta$. Then, given the definition of the $M$ matrix in the main text, the fermionic force can be straightforwardly computed as
$F_{i,\tau}$ $=X^{\dagger}\left[M^{\dagger}(\phi)\left(\frac{\partial}{\partial\phi_{i,\tau}}M(\phi)\right)+\left(\frac{\partial}{\partial\phi_{i,\tau}}M^{\dagger}(\phi)\right)M(\phi)\right]X$ (S3)
$=\sum_{\tau^{\prime}=1}^{N_{\tau}-1}\left[X_{\tau^{\prime}}^{\dagger}\left(B_{{\tau^{\prime}}}^{\dagger}\frac{\partial B_{{\tau^{\prime}}}}{\partial\phi_{i,\tau}}\right)^{\dagger}X_{\tau^{\prime}}+X_{\tau^{\prime}}^{\dagger}\left(B_{{\tau^{\prime}}}^{\dagger}\frac{\partial B_{{\tau^{\prime}}}}{\partial\phi_{i,\tau}}\right)X_{\tau^{\prime}}+X_{{\tau^{\prime}}+1}^{\dagger}\left(-\frac{\partial B_{{\tau^{\prime}}}}{\partial\phi_{i,\tau}}\right)X_{{\tau^{\prime}}}+X_{{\tau^{\prime}}}^{\dagger}\left(-\frac{\partial B_{{\tau^{\prime}}}}{\partial\phi_{i,\tau}}\right)^{\dagger}X_{{\tau^{\prime}}+1}\right]$
$+X_{N_{\tau}}^{\dagger}\left(B_{{N_{\tau}}}^{\dagger}\frac{\partial B_{{N_{\tau}}}}{\partial\phi_{i,\tau}}\right)^{\dagger}X_{N_{\tau}}+X_{N_{\tau}}^{\dagger}\left(B_{{N_{\tau}}}^{\dagger}\frac{\partial B_{{N_{\tau}}}}{\partial\phi_{i,\tau}}\right)X_{N_{\tau}}+X_{1}^{\dagger}\left(\frac{\partial B_{1}}{\partial\phi_{i,{N_{\tau}}}}\right)X_{{N_{\tau}}}+X_{{N_{\tau}}}^{\dagger}\left(\frac{\partial B_{1}}{\partial\phi_{i,{N_{\tau}}}}\right)^{\dagger}X_{1}.$ (S4)
The remaining step is to evaluate $\frac{\partial B_{\tau^{\prime}}}{\partial\phi_{i,\tau}}$. For symmetric checkerboard decomposition of the form $B_{\tau}=\prod_{n}e^{V_{\tau,n}/2}$ as shown in the main text, its derivative with respect to $\phi_{i}$ is
$\frac{\partial B_{\tau^{\prime}}}{\partial\phi_{i,\tau}}$ $=\delta_{\tau\tau^{\prime}}\sum_{m}\prod_{l<m}e^{V_{\tau,l}/2}\frac{\partial e^{V_{\tau,m}/2}}{\partial\phi_{i,\tau}}\prod_{n>m}e^{V_{\tau,n}/2}.$ (S5)
Thus, each term in Eq. (S4) can be expressed as a series of matrix-vector multiplications, where the matrix is of the form $e^{V_{\tau,n}}$. Each matrix-vector multiplication has a time complexity of $O(N_{\tau}V_{s})$. So, the total time complexity of computing the fermionic force is $O(N_{\tau}V_{s})$.
Figure S1: The average sign of $\det M$ as a function of $J$ for various lattice sizes shown as $N_{\tau}\times L^{2}$ computed for the non-compact QED3with $K=0$, by DQMC method. For $J\lesssim 2.0$, the $\det M$ are almost all positive. Thus, the ergodicity issue of HQMC is not severe in this regime of $J$.
Figure S2: The spin structure factor, which is Fourier transformation of spin-spin correlation function, evaluated at $X=(\pi,\pi)$ point, as a function of $J$, for various lattice sizes $N_{\tau}\times L^{2}$, computed for the non-compact QED3 with $K=0$. The plot compares the results between HQMC and DQMC. For $J\lesssim 2.0$, the spin structure factor from HQMC agrees well with that from DQMC; they only deviate from each other at large $J$, which is attributed to the ergodicity issue of the HQMC algorithm. This shows that in the compact QED3 coupled with fermions, the ergodicity issue of HQMC is negligible for small $J$ deep in the putative $U(1)$ DSL phase.
Figure S3: The average sign of $\det M$ as a function of $J$ for various lattice sizes shown as $N_{\tau}\times L^{2}$ computed for the compact QED3with $K=1$, by DQMC method. For $J\lesssim 2.2$, the $\det M$ are almost all positive. Thus, the ergodicity issue of HQMC is not severe in this regime of $J$.
Figure S4: The spin structure factor, which is Fourier transformation of spin-spin correlation function, evaluated at $X=(\pi,\pi)$ point, as a function of $J$, for various lattice sizes $N_{\tau}\times L^{2}$, computed for the compact theory with $K=1$. The plot compares the results between HQMC and DQMC. For $J\lesssim 2.2$, the spin structure factor from HQMC agrees well with that from DQMC; they only deviate from each other at large $J$, which is attributed to the ergodicity issue of the HQMC algorithm. This shows that in the compact QED3 coupled with fermions, the ergodicity issue of HQMC is negligible for small $J$ deep in the putative $U(1)$ DSL phase.
SII. Benchmark data
As mentioned in the main text, the HQMC algorithms applied to the QED3 has an ergodicity issue, in the sense that the negative $\det M$ region cannot be reached going from positive $\det M$ region if only HQMC updates are proposed. However, we find that if $J$ is small enough, the negative $\det M$ region will have a negligible possibility of being sampled, thus the HQMC method is valid for such values of $J$. In this section, we use the DQMC algorithm, which is ergodic, to benchmark HQMC and find the region of $J$ where the HQMC result aegrees with the DQMC result. We will show that the $J=1.25$ we chose throughout the paper is inside such valid region.
Fig. S1 shows the averaged sign of $\det M$ as a function of $J$ computed by DQMC for the non-compact QED3. It shows that, at large $J$ where the gauge field is highly dynamic (great fluctuation along $\tau$ direction), the negative $\det M$ samples are important and lower the average sign to be <1.0<1.0. But as $J$ decreases to below $2.0$ where the gauge field becomes more static, the sign is almost $1.0$. Thus, the negative $\det M$ samples are negligible.
Next, we compare the HQMC result with the DQMC result for various $J$. Fig. S2 shows the spin structure factor, a Fourier transformation of the spin-spin correlation function, evaluated at $X=(\pi,\pi)$ point. This observable serves as a detect of the antiferromagnetic (AFM) order. In this figure, in the large $J$ region, HQMC result deviates from the DQMC result as expected since the negative $\det M$ gauge configurations are important and not accessible to HQMC updates. But for $J\lesssim 2.0$, they agree with each other, indicating the ergodicity issue is negligible for these $J$.
Benchmarking HQMC with DQMC on compact QED3 is similar to that on non-compact QED3  which are shown in Fig. S3 and Fig. S4. Again, in the small $J$ region of around $J\lesssim 2.2$, the signs of the $\det M$ are almost always positive, the AFM order parameters between HQMC and DQMC agree with each other, and thus the ergodicity issue of HQMC is negligible.
Figure S5: The log plot of the spin-spin correlation as a function of distance $r$ for compact QED3with $K=1$, at $J=1.25$, for various lattice sizes shown as $N_{\tau}\times L^{2}$ in the legend. To avoid the even-odd oscillation in the finite-size data, we only plot the data of the distance $r=\textmd{odd}$ points. The gray square, triangular and diamond dots show the results computed from the DQMC to benchmark the HQMC result of the corresponding lattice sizes. At the large distance and the large system sizes, the data are manually fitted by the black line, which displays a power law decay with scaling dimension of around $3.2$.
Figure S6: The log-log plot of the bond-bond correlation (along $\hat{x}$ direction) as a function of distance $r$ for compact QED3with $K=0$, at $J=1.25$, for various lattice sizes shown as $N_{\tau}\times L^{2}$ in the legend. Notice that $C_{S}^{\uparrow\downarrow}=C_{S}^{\downarrow\uparrow}$, so we only need to measure $C_{S}^{\uparrow\downarrow}$ here. To avoid the even-odd oscillation in the finite-size data, we only plot the data of the distance $r=\textmd{odd}$ points. At the large distance and the large system sizes, the data are manually fitted by the black line, which displays a power law decay with scaling dimension of around $3.2$.
Figure S7: The log-log plot of the flux-flux correlation as a function of imaginary time $\tau$ for compact QED3  with $K=1$ and $J=1.25$, for various lattice sizes, shown as $N_{\tau}\times L^{2}$ in the legend. At the large distance and the large system sizes, the data are manually fitted by the black line, which displays a power law decay with scaling dimension of around $4.1$. As lattice sizes increases, the slope still shows finite-size oscillation behavior up to size $N_{\tau}\times L^{2}=660\times 66^{2}$.
SIII. Data for the compact QED3 model
In this section, we present the results of the spin, bond and flux correlation for the compact QED3 model, which is parallel to the results of noncompact QED3 in the main text. We will again analyze the scaling behavior of these observables and demonstrate how they reveal the conformal field information in the $U(1)$ DSL critical phase.
Fig. S5 shows the log plot of the spin-spin correlation as a function of lattice distance $r$ for compact QED3 at $J=1.25$, for various lattice sizes. The noise level is at around $7\times 10^{-6}$. As lattice size increases, the log plot of the spin-spin correlation function consistently displays a linear decay behavior at large distance $r$, until it reaches the noise level. This agrees with the power law decay behavior of the fermionic observable, which manifests the property of the $U(1)$ DSL critical phase. The linear decay is manually fitted by the black solid line which gives a power of around $3.2$. This agrees with the scaling dimension of fermion bilinear operators $2\Delta_{\rm adj}\in(2.8,3.4)$ predicted by the CFT, where the fermion bilinear operators serve as the adjoint representation of the emergent $SU(4)$ symmetry, as introduced in the main text and Refs. [3, 18]. In Fig. S5, we also plot the DQMC results for several small lattice sizes to benchmark the HQMC results, and they agree well with each other.
Similarly, in Fig. S6 which shows the log plot of the bond-bond correlation as a function of distance $r$, the data points for various lattice sizes also consistently converge to a straight line, until they reach a noise level of $2\times 10^{-5}$. The points are large distance $r$ are manually fitted with the black solid line, which gives the power of around $3.2$. This again agrees with the scaling dimension of fermion bilinear operators $2\Delta_{\rm adj}\in(2.8,3.4)$ predicted by the CFT.
Fig. S7 shows the log plot of the flux-flux correlation as a function of the imaginary time $\tau$ for compact QED3 model with $K=1$ and $J=1.25$. As lattice size increases, the flux-flux correlation at large $\tau$ consistently converge to a clear linear behavior before they reach a noise level of around $2\times 10^{-3}$. Such limit behavior is again manually fitted by the black solid line, which gives a power of around $4.1$, and is distinct from the dashed line with the power of $3.0$. This verifies the fact that the scaling dimension of the conserved current in $U(1)$ DSL phase is $\Delta_{J}=2$.
In a nutshell, the results of spin, bond and magnetic flux correlation function of compact QED3 does not have an essential difference from the results of noncompact QED3 shown in the main text. This shows that, even though in theory the monopole operator in compact QED3 may create relevant perturbation, gap out the gauge field and destroy the $U(1)$ DSL phase at any $J$, in actual numerical simulation, such signature is not seen at a finite lattice size of up to $N_{\tau}\times L^{2}=660\times 66^{2}$ at $J=1.25$.
SIV. Derivation of fermionic correlation functions
In this section we give a detailed derivation of the fermionic correlation functions measured in HQMC, which includes the spin correlation function and bond correlation function.
As introduced in the main text, the spin operator for $N_{f}=2$ is defined as
$S^{\alpha\beta}(i)$ $=\psi_{i\alpha}^{\dagger}\psi_{i\beta}-\frac{1}{2}\delta_{\alpha\beta}\sum_{\gamma}\psi_{i\gamma}^{\dagger}\psi_{i\gamma}.$ (S6)
The bond operator along the nearest-neighbor bond in $\hat{x}$ direction is defined as
$B_{i}$ $=\sum_{\alpha}(\psi^{\dagger}_{i,\alpha}e^{i\phi_{ij}}\psi_{i+\hat{x},\alpha}+\text{h.c.}),$ (S7)
and $\alpha,\beta\in\left\{\uparrow,\downarrow\right\}$ are spin flavor indices. Then, the spin correlation function is
$C_{\textmd{S}}^{\uparrow\downarrow}(i,j)$ $=\langle S^{\uparrow\downarrow}(i)S^{\downarrow\uparrow}(j)\rangle_{\psi\phi}$ (S8)
$=\langle\psi^{\dagger}_{i\uparrow}\psi_{i\downarrow}\psi^{\dagger}_{j\downarrow}\psi_{j\uparrow}\rangle_{\psi\phi}$ (S9)
$=\langle\bar{G}_{\uparrow}(i,j)G_{\downarrow}(i,j)\rangle_{\phi},$ (S10)
where the last row utilizes Wick’s theorem, and the subscripts $\psi$ and $\phi$ mean the average over the fermionic field and the bosonic gauge field, respectively. The equal-time Green’s functions are defined as
$G_{\alpha}(i,j)$ $=\langle\psi_{i\alpha}\psi^{\dagger}_{j\alpha}\rangle_{\psi},$ (S11)
$\bar{G}_{\alpha}(i,j)$ $=\langle\psi^{\dagger}_{i\alpha}\psi_{j\alpha}\rangle_{\psi}$ (S12)
$=\delta_{ij}-G_{\alpha}(j,i).$ (S13)
In this problem, the Green’s function of the two fermion flavors are equal $G_{\uparrow}=G_{\downarrow}$. So we will omit the flavor index in the Green’s functions next. Similarly, the bond correlation function is
$C_{\textmd{B}}(i,j)$ $=\langle B_{i}B_{j}\rangle_{\psi\phi}-\langle B_{i}\rangle_{\psi\phi}\langle B_{j}\rangle_{\psi\phi}.$ (S14)
Here,
$\langle B_{i}B_{j}\rangle_{\psi\phi}=2$ $\langle\bar{G}(i,j+\hat{x})G(i+\hat{x},j)e^{i(\phi_{i,i+\hat{x}}+\phi_{j,j+\hat{x}})}$
$+$ $\bar{G}(i,j)G(i+\hat{x},j+\hat{x})e^{i(\phi_{i,i+\hat{x}}-\phi_{j,j+\hat{x}})}$
$+$ $\bar{G}(i+\hat{x},j+\hat{x})G(i,j)e^{i(-\phi_{i,i+\hat{x}}+\phi_{j,j+\hat{x}})}$
$+$ $\bar{G}(i+\hat{x},j)G(i,j+\hat{x})e^{-i(\phi_{i,i+\hat{x}}+\phi_{j,j+\hat{x}})}\rangle_{\phi}$
$+4\langle$ $\left[\bar{G}(i,i+\hat{x})e^{i\phi_{i,i+\hat{x}}}+\bar{G}(i+\hat{x},i)e^{-i\phi_{i,i+\hat{x}}}\right$]
$\left[\bar{G}(j,j+\hat{x})e^{i\phi_{j,j+\hat{x}}}+\bar{G}(j+\hat{x},j)e^{-i\phi_{j,j+\hat{x}}}\right]\rangle_{\phi},$ (S15)
where the coefficients come from the summation of the fermionic flavors.
$\langle B_{i}\rangle_{\psi\phi}\langle B_{j}\rangle_{\psi\phi}=4\langle$ $\bar{G}(i,i+\hat{x})e^{i\phi_{i,i+\hat{x}}}+\bar{G}(i+\hat{x},i)e^{-i\phi_{i,i+\hat{x}}}\rangle_{\phi}$
$\langle\bar{G}(j,j+\hat{x})e^{i\phi_{j,j+\hat{x}}}+\bar{G}(j+\hat{x},j)e^{-i\phi_{j,j+\hat{x}}}\rangle_{\phi},$ (S16)
We denote the vector $b_{i}=\bar{G}(i,i+\hat{x})e^{i\phi_{i,i+\hat{x}}}+\bar{G}(i+\hat{x},i)e^{-i\phi_{i,i+\hat{x}}}$, then
$C_{\textmd{B}}(i,j)=2$ $\langle\bar{G}(i,j+\hat{x})G(i+\hat{x},j)e^{i(\phi_{i,i+\hat{x}}+\phi_{j,j+\hat{x}})}$
$+$ $\bar{G}(i,j)G(i+\hat{x},j+\hat{x})e^{i(\phi_{i,i+\hat{x}}-\phi_{j,j+\hat{x}})}$
$+$ $\bar{G}(i+\hat{x},j+\hat{x})G(i,j)e^{i(-\phi_{i,i+\hat{x}}+\phi_{j,j+\hat{x}})}$
$+$ $\bar{G}(i+\hat{x},j)G(i,j+\hat{x})e^{-i(\phi_{i,i+\hat{x}}+\phi_{j,j+\hat{x}})}\rangle_{\phi}$
$+$ $4\left(\langle b_{i}b_{j}\rangle_{\phi}-\langle b_{i}\rangle_{\phi}\langle b_{j}\rangle_{\phi}\right),$ (S17)
where the first term is the intrinsic fermionic correlation while the second term is the correlation from the gauge field fluctuation.